OpenStudy (anonymous):

Clyde has $900 invested at 8%. How much money must Clyde invest at 3% so that he will assume a 6% return altogether on the two investments? I want to solve in a table

OpenStudy (tkhunny):

T = The Amount to invest at 3% D = The amount of interest earned by T. 900 * 0.08 = 72 T * 0.03 = D (900 + T) * 0.06 = (72 + D) Now what?

OpenStudy (anonymous):

now what do i do

OpenStudy (tkhunny):

Solve? That was the table.